We have a 6% gallon jug of bleach and we want to make a 1% solution of bleach, how much water is needed to make this 1% batch? 1 gal. will make a 3% solution of bleach and we now have 2 gals so then if we added 2 more gals we would have a 1.5% solution with 4 gals, 3 of which were pure water added. Just looking for a good formula to get the right answer, came up with this but not sure if I am correct .06(1Gal) + 0(x)= .01(1+x) so .06 + 0= .01+.01x
Answer by stanbon(75887)

We have a 6% gallon jug of bleach and we want to make a 1% solution of bleach, how much water is needed to make this 1% batch? 1 gal. will make a 3% solution of bleach and we now have 2 gals so then if we added 2 more gals we would have a 1.5% solution with 4 gals, 3 of which were pure water added. Just looking for a good formula to get the right answer, came up with this but not sure if I am correct .06(1Gal) + 0(x)= .01(1+x) so .06 + 0= .01+.01x
--------------------
That equation looks good.
What you have is
bleach + bleach = bleach
0.06*1 + 0*x = 0.01(1+x)
Multiply thru by 100 to get:
6 + 0 = 1+x
x = 5 gallon (amt of water you need to add to get a 1% solution)
==================
Cheers,
Stan H.
==================